A parallel-plate capacitor has plates with
area of cross-section 100 cm² and separated
by 1.0 cm. A potential difference of 100 V
is applied when no dielectric is present. A
dielectric slab of relative permittivity 5 and
thickness 0.5 cm is introduced between the
plates. Calculate the free charge on the
capacitor plates.
1
Expert's answer
2019-05-28T08:59:18-0400

According to the formula of capacity for parallel-plate capacitor:


"C=(\epsilon*\epsilon_0*S)/d (1)"

Where

"\epsilon" - relative permittivity

"\epsilon_0" - electric constant, "\epsilon_0=8.854*10^{−12} F⋅m^{−1}"

"S" - area of plates

"d" - distance between plates


After introducing the slab between the plates, the system could be described as 3 capasitors, connected together in series.

First capasitor capacity, according to (1):


"C_1=(\epsilon*\epsilon_0*S)/x_1 (2)"

Where

"x_1" - distance from first plate to slab.


Second capasitor capacity, according to (1):


"C_2=(\epsilon_2*\epsilon_0*S)/d_2 (3)"

Where

"\epsilon_2" - relative permittivity of a slab.

"d_2" - thickness of a slab


Third capasitor capacity, according to (1):

"C_3=(\epsilon*\epsilon_0*S)/x_2 (4)"

Where

"x_2" - distance from last plate to slab.

According the formula of connectected in deries capacitors:


"1\/C=1\/C_1+1\/C_2+1\/C_3 (5)"

Using (2), (3) and (4):


"1/C=x_1/(\epsilon*\epsilon_0*S)+d_2/(\epsilon_2*\epsilon_0*S)+x_2/(\epsilon*\epsilon_0*S) (6)"

"\implies"

"C=(\epsilon*\epsilon_2*\epsilon_0*S)/(\epsilon_2*(x_1+x_2)+\epsilon*d_2) (7)"

The distance between the plates do not changes "\implies" "x_1+x_2=d-d_2 (8)"

Where

"d" - sictance between plates

Using (7), (8):


"C=(\epsilon*\epsilon_2*\epsilon_0*S)/(\epsilon_2*(d-d_2)+\epsilon*d_2) (9)"

Acording the definition of capabitor capacity:


"C=Q/U \implies Q=C*U (10)"

Where

"Q" - charge on the plates

"U" - potential difference

Using (9), (10)


"Q=U*(\epsilon*\epsilon_2*\epsilon_0*S)/(\epsilon_2*(d-d_2)+\epsilon*d_2)"

Using the numbers in SI:

"U=100V"

"\epsilon=1"

"\epsilon_0=8.854*10^{−12} F⋅m^{−1}"

"\epsilon_2=5"

"S=100cm^2=10^{-2}m^2"

"d=1cm=10^{-2}m"

"d_2=0.5cm=0.5*10^{-2}m"


Answer: the free charge on the capacitor plates is "Q=1.61*10^{-9} \u0421"